OpenStudy (anonymous):

Hey There Everyone! :) I need some help. :) Find a third-degree polynomial function such that f(0) = -24 and whose zeros are -1, 2, and -3. Using complete sentences, explain how you found it.

OpenStudy (campbell_st):

well start with the zeros if x = -1 is a zero then (x + 1) is a factor x = 2 is a zero then (x -2) is a factor x = -3 is a zero then (x + 3 is a factor so the polynomial is f(x) =a(x +1)(x+3)(x-2) to find the value of a, substitute x = 0 and f(0) = 24 then evaluate last step with be writing the polynomial in expanded form. hope it helps

OpenStudy (anonymous):

You are my hero. :)

OpenStudy (campbell_st):

lol... thanks, glad to help